The test scores of 40 students are summarized in the frequency table below. Find the standard deviation.
Score Students
50-59 12
60-69 8
70-79 10
80-89 5
90-99 5
Ive never done standard deviation on numbers that are ranged, so I dont even know how to set up the problem.

Answer by stanbon(75887)

The test scores of 40 students are summarized in the frequency table below. Find the standard deviation.
Score Students
50-59 12
60-69 8
70-79 10
80-89 5
90-99 5
Ive never done standard deviation on numbers that are ranged, so I dont even know how to set up the problem.
------------------
You use the class mark as the representative value for each class.
The class mark for the 1st class is (59+50)/2 = 109/2 = 54.4
And the frequency of that value is 12.
---------------
Do the same for each class. Then determine the std.
========================================================
Cheers,
Stan H.